# Lift-Attend-Splat: Bird’s-eye-view camera-lidar fusion using transformers

### v1

@@ 5 Experiments
We use Dual-Swin-Tiny [27] with a feature pyramid network [29] and VoxelNet [65] as our camera and lidar encoders respectively. For our object detection head, we use the transformer-decoder-based module from TransFusion-L [1]. We use our Lift-Attend-Splat method, described in Sec. 4, to project camera features into BEV space. We then fuse camera and lidar features using simple concatenation and convolution. Following [33] the RPN part of VoxelNet is applied to the merged feature. We ablate alternative choices for the fusion architecture in Sec. 5.4.
@@
We use an image resolution of 800x448 for inputs to the camera encoder, which is downsampled 8x by the camera encoder into per-camera feature maps of shape 100x56. For VoxelNet, we follow the settings of [28]. During training, we set a maximum of 90k non-empty voxels which we increase to 180k during inference. We use a lidar BEV grid centred around the ego with dimensions $108\text{m}\times 108\text{m}$ and 0.075m cell size. This is downsampled 8x by the lidar encoder to the $180\times 180$ grid into which the camera features are projected. We construct the intermediate projected horizon with 143 uniformly spaced depth bins ranging from 1m to 72m. For the projection of camera features into BEV, we use the original transformer [48] as our encoder-decoder architecture, with one encoder layer, one decoder layer, $d_{\text{model}}=256$ and $d_{\text{ff}}=512$. We replace the ReLU [36] activation function with GeLU [13], use learnable position embeddings [10] in place of sinusoidal encodings and normalise features before each sublayer [56]. We tie the parameters in each of the 8 heads of our attention blocks. For the object detection head, we use 200 and 300 queries during training and inference respectively.
@@
Our lidar backbone is pretrained on 8 GPUs with batch-size of 1/GPU following the schedule presented in [1], with CBGS [66] and copy-paste augmentation [58]. We initialise the camera backbone with weights pretrained on nuImages [2]. We freeze the lidar backbone and train the camera backbone, projection, fusion and detection head together for 20 epochs using 8 GPUs with batch-size of 4/GPU. We use the AdamW optimiser [34] with a maximum learning rate of $5\times 10^{-5}$ for the camera backbone and $1\times 10^{-3}$ for all other components. We apply the following augmentations: random mirror flips in Y dimensions, affine transformations (rotations and scale), and camera-lidar copy-paste augmentations [50].
@@
We present results for the task of 3D object detection on Tab. 1. Compared to baselines based on the Lift-Splat projection [28, 33], our method shows improvements both on the validation and test splits of the nuScenes dataset. In particular, we show substantial improvements in both mAP ($+1.1$) and NDS ($+0.4$) on the test split. Since the lidar backbone is frozen and similar in all approaches, this shows that our model is better able to leverage camera features. Per-class object detection results as well as comparison to other methods can be found in Tab. S2 and Sec. B.1.
We can analyse further the performance of our model by clustering objects together depending on their distances from the ego and on their sizes, see Fig. 3. We see that the bulk of the improvements comes from objects located at large distances and of small sizes. These are situations for which monocular depth estimation is particularly difficult which could explain why our model fares better in these cases. Note that even though far-away and small objects contain less lidar points, our model is still able to leverage camera features effectively even though the context given by the lidar is weaker.
Finally, we show results that use test-time-augmentations (TTA) and model ensembling at the bottom of Tab. 1. We perform TTA over a combination of mirror and rotation augmentations and ensemble models with cell resolutions of 0.05m, 0.075m and 0.10m. We first apply TTA at each cell resolution and then merge the resulting boxes using Weighted Boxes Fusion (WBF) [47]. Unsurprisingly, our method shows excellent scaling with respect to these techniques and outperforms BEVFusion [33] on the nuScenes validation set. More details can be found in Sec. B.3.
@@
${}^{\ddagger}$ [33] 73.7 74.9 75.0 76.1
${}^{\ddagger}$ 74.6 75.1 - -
Ours w/ ${}^{\ddagger}$ 75.7 76.0 75.5 74.9
Table 1: Our method compared to LiftSplat-based camera-lidar fusion baselines on the validation and test splits of the nuScenes dataset. TFA: Temporal Feature Aggregation. ${}^{\ddagger}$ denotes test-time augmentations and model ensembling.
@@ 5.2 Qualitative analysis
OursRefer to caption (a) BEVFusion [33]Refer to caption (b)
OursRefer to caption BEVFusion [33]Refer to caption Refer to caption Refer to caption Refer to caption (c)
Camera+lidar Camera only
@@
Figure 4: (a, b) Visualisation of where camera features of ground-truth objects are projected onto the BEV grid for our method compared to BEVFusion [33]. We observe that our method is able to place camera features around objects more narrowly than BEVFusion, which is based on monocular depth estimation. (c) Comparison of object saliency maps for models trained with camera only (top) and camera and lidar (bottom). When trained with both camera and lidar, our model selects camera features in an area that is different than when trained with camera only, while [33] behaves similarly in both settings. Saliency maps and original images show the corresponding ground-truth box for context.
@@
We visualise where camera features are projected onto the BEV grid and compare our method to BEVFusion [33]. For our method, we examine the attention map of the final cross-attention block in the transformer, averaged over all attention heads. For BEVFusion, we use the the monocular depth estimate to establish the strength of correspondence between positions in camera and BEV space. We consider only the pixels corresponding to ground-truth objects when calculating the total weight of projected camera features in BEV, see Sec. B.2 for details. As can be seen in Fig. 3(a) (left), our method places camera features predominantly in regions where ground-truth bounding boxes are present. This indicates that it can effectively leverage the lidar point cloud as a context in order to project camera features at the relevant location in BEV. Compared to BEVFusion shown in Fig. 3(b), the distribution of features appears more narrowly localised and stronger around objects. This could be due to the fact our projection mechanism does not require the weights of the camera features to be normalised along their ray, giving our model more flexibility to place features at the desired location. It is interesting to note that even though our method also projects camera features outside of ground-truth boxes in BEV, the strength of the activation in these regions is suppressed by the fusion module, see Fig. S4. This is consistent with our findings in Sec. 3, where we showed that the latter part of the model is able to effectively suppress camera features at the wrong location. More examples can be found in Sec. B.2.
We further explore which pixels in the camera images are most attended to using saliency maps [46]. These are derived by computing the gradient of the maximum class logit with respect to a camera image $I_{j}$, given object query index $i$ and probabilities $z$, as:
$\left.\frac{\partial z_{i,\hat{c}}}{\partial I}\right|_{I_{j}}\text{where }\hat{c}=\arg\max_{c}z_{i,c}.$ (6)
@@
They allow us to visualise the contribution of individual pixels to the final prediction for a selected object, see Fig. S2 (right). Interestingly, we observe that when trained with both camera and lidar, our model tends to select camera features at different locations than when trained with cameras only. In the absence of lidar, our method selects camera features across the entirety of the object, while in the presence of both lidar and cameras, the model selects camera features mainly from the upper part of the object. We observe that this pattern is mostly prevalent for objects close to the ego which are well-represented by lidar point clouds but fades away for far-away objects or objects with few lidar points such as pedestrians, see Fig. S2 for more examples. We hypothesise that our projection architecture enables the model to select camera features that complement best the information encoded in lidar, resulting in differing attention patterns between camera-only and fusion settings. This pattern is less present in BEVFusion [33], which attends to the broader neighbourhood of pixels surrounding the selected object in both cases.
@@ 5.3 Temporal feature aggregation
We train TFA models on sequences of 3 frames for 10 epochs starting from a pretrained object detection head, lidar and camera backbones (from our single frame model). During training, the lidar and camera backbones are kept frozen. For inference, we accumulate BEV features for the entire length of a run, yielding detections at each time step. Table 1 shows that temporal feature aggregation boosts object detection performance significantly, both without and with TTA and model ensembling.
@@ 5.4 Ablation experiments
Fusion module:
 Cat+Conv 70.43 71.9
@@
Number of decoder $*{}^{*}$:
 1 block 70.29 71.9
@@
Table 2: Impact of various modifications of our model on 3D object detection performance: (i) feature fusion module, (ii) number of transformer decoder blocks in the “Attend” stage, (iii) number of frames in Temporal Feature Aggregation (TFA). $*{}^{*}$ indicates frozen camera backbone.
We ablate some design choices for our method and show their impact on object detection performance on Tab. 2. For all the ablation experiments, we use a simpler training schedule with 10 epochs, batch accumulation instead of full batch training and no camera augmentations. We first analyse the impact of different implementations of the fusion module: we compare a simple skip connection (add), a small concatenation and convolution layer (Cat+Conv as in [33]) and a gated sigmoid block [28]. We find all of them to perform very similarly, with Cat+Conv performing slightly beter with respect to mAP, contrary to findings of [28]. We also ablate the number of transformer decoder blocks in the “Attend” stage of our projection and show that increasing their number does lead to small improvement in mAP. This suggests that our method can scale up with increased compute but we use a single decoder block in our experiments as it provides good balance between quality and performance. Finally, we also see good improvement in detection score when increasing number of frames in Temporal Feature Aggregation during training.
@@ 6 Conclusion
In this work, we have analysed the role of monocular depth prediction in recent state-of-the-art camera-lidar fusion methods and showed that, surprisingly, improvements in depth estimation did not lead to better object detection performance. Strikingly, we also showed that removing depth estimation altogether did not worsen performance significantly. This led us to hypothesise that relying on monocular depth estimation could be an unnecessary architectural bottleneck when fusing camera and lidar, and prompted us to introduce a novel fusion method that directly combines camera and lidar features using a simple attention mechanism. Compared to projecting camera features using monocular depth, our method allows camera features to contribute to multiple locations in BEV space and gives our model greater flexibility to select complementary camera and lidar features. Finally, we validated the effectiveness of our method on the nuScenes dataset and showed that it improves object detection performance over baselines based on monocular depth estimation and showcased the role of attention as a key contributor to these improvements. We hope that our work will motivate discussions around the role of monocular depth prediction in camera-lidar fusion and motivate further developments in multi-modal perception more generally.
@@ Acknowledgements
We thank our FiveAI and Bosch colleagues, especially Tom Joy and Anthony Knittel, for their valuable feedback and suggestions on the manuscript. We are grateful to Bhavesh Garg for his initial experiments into depth supervision, whilst interning at FiveAI, which helped spawn the basis of this work. Finally, we thank Blaine Rogers for the early work on fusion at FiveAI from which this work developed.
@@ A.4 Extended depth quality results
We evaluate the performance of the depth classifier using five of the metrics proposed in [9]: root mean squared error (RMSE), root mean squared logarithmic error (RMSLE), mean absolute relative error (Abs. Rel.), mean squared relative error (Sq. Rel.) and fraction outside 125% (Frac. 125). We show all the metrics for 2 different methods of translating the classification output into a depth map: “mode” — in which we use the bin with maximum probability, and “mean” — where the predicted depth is the weighted average of all the bins. We take the average of these quantities over all predictions made by the depth classifier for which we have ground truth to compute the metrics. Camera feature cells for which lidar depth is unspecified are ignored. We compare BEVFusion and four different variants: adding depth supervision using Eq. 2 with various weights $\lambda$, using lidar depth maps instead of monocular depth estimation (lidar), using a pretrained and frozen depth classifier (pretrained), and finally removing depth estimation altogether by projecting camera features at all depths uniformly using Eq. 3 (no depth). Quantitative results can be seen in Tab. S1 and qualitative comparisons in Fig. S1.
@@ B Detailed experimental results
Table S2: Per-class object detection scores on the nuScenes validation set (top) and test set (bottom). TFA: Temporal Feature Aggregation. ${}^{\ddagger}$ indicates ensembling + TTA.
@@ B.2 Detailed qualitative results
OursRefer to caption BEVFusion [33]Refer to caption Refer to caption Refer to caption Refer to caption (a)
OursRefer to caption BEVFusion [33]Refer to caption Refer to caption Refer to caption Refer to caption (b)
Cam+lidar Cam-only
OursRefer to caption BEVFusion [33]Refer to caption Refer to caption Refer to caption Refer to caption (c)
OursRefer to caption BEVFusion [33]Refer to caption Refer to caption Refer to caption Refer to caption (d)
Cam+lidar Cam-only

### v3

@@ 5 Experiments
We use Dual-Swin-Tiny [30] with a feature pyramid network [32] and VoxelNet [70] as our camera and lidar encoders respectively. Our object detection head is the transformer-decoder-based module from TransFusion-L [1]. We use our Lift-Attend-Splat method, described in Sec. 4, to project camera features into BEV space. We then fuse camera and lidar features using simple concatenation and convolution. Following [36] the RPN part of VoxelNet is applied to the merged feature. We ablate alternative choices for the fusion architecture in Sec. 5.4.
@@
Inputs to the camera encoder have resolution 800x448, which it downsamples by 8x into per-camera feature maps of shape 100x56. For VoxelNet, we follow the settings of [31]. We set a maximum of 90k non-empty voxels during training, increased to 180k for inference. We use an ego-centric BEV grid with dimensions $108\text{m}\times 108\text{m}$ and 0.075m cell size. This is downsampled 8x by the lidar encoder to the $180\times 180$ grid into which the camera features are projected. We construct the intermediate projected horizon with 143 uniformly spaced depth bins ranging from 1m to 72m. For the projection of camera features into BEV, we use the original transformer [51] as our encoder-decoder architecture, with one encoder layer, one decoder layer, $d_{\text{model}}=256$ and $d_{\text{ff}}=512$. We replace the ReLU [39] activation function with GeLU [15], use learnable position embeddings [11] in place of sinusoidal encodings and normalise features before each sublayer [60]. We tie the parameters in each of the 8 heads of our attention blocks. For the object detection head, we use 200 and 300 queries during training and inference respectively.
@@
Our lidar backbone is pretrained on 8 GPUs with batch-size of 1/GPU following the schedule presented in [1], with CBGS [71] and copy-paste augmentation [63]. We initialise the camera backbone with weights pretrained on nuImages [2] by [31]. We freeze the lidar backbone and train the camera backbone, projection, fusion and detection head for 20 epochs using 8 GPUs with a batch size of 4/GPU. We use the AdamW optimiser [37] with a maximum learning rate of $5\times 10^{-5}$ for the camera backbone and $1\times 10^{-3}$ for all other components. We apply the following augmentations: mirroring in the Y dimension, global rotation and scale, and camera-lidar copy-paste [53].
@@
We present results for the task of 3D object detection on Tab. 1. Compared to baselines based on the Lift-Splat projection [31, 36], our method shows improvements on the validation and test splits of the nuScenes dataset. In particular, we show substantial improvements in both mAP ($+1.1$) and NDS ($+0.4$) on the test split. The lidar backbone is frozen and similar in those approaches, showing that our model is better able to leverage camera features. Our method outperforms the more recent fusion algorithms TransFusion, DeepInteraction, and FUTR3D by a substantial margin, performs similarly to UniTR and MSMDFusion, and slightly underperforms CMT. These results show that our simple modification of BEVFusion is successful in raising its performance on par with recent SOTA methods. We leave extending our method with multi-scale feature fusion (used by MSMDFusion) and query denoising (used by CMT) to future work. Per-class object detection results and comparisons can be found in Tab. S2 and Sec. B.1.
We can analyse further the performance of our model by clustering objects together depending on their distances from the ego and on their sizes, see Fig. 4. We see that the bulk of the improvements comes from objects located at large distances and of small sizes. These are situations for which monocular depth estimation is particularly difficult which could explain why our model fares better in these cases. Note that even though far-away and small objects contain fewer lidar points, our model is still able to leverage camera features effectively even though the context given by the lidar is weaker.
We show results that use test-time-augmentations (TTA) and model ensembling at the bottom of Tab. 1. We perform TTA over a combination of mirror and rotation augmentations and ensemble models with cell resolutions of 0.05m, 0.075m and 0.10m. We first apply TTA at each cell resolution and then merge the resulting boxes using Weighted Boxes Fusion (WBF) [50]. Unsurprisingly, our method shows excellent scaling for these techniques and outperforms BEVFusion [36] on the nuScenes validation set. More details can be found in Sec. B.3.
@@
BEVFusion [36] 73.7 74.9 75.0 76.1
Ours 74.6 75.1 - -
Ours w/ TFA 75.7 76.0 75.5 74.9
Table 1: Object detection performance on the validation and test splits of the nuScenes dataset. TFA: Temporal Feature Aggregation. denotes test-time augmentations and model ensembling.
Refer to captionRefer to caption
@@ 5.2 Qualitative analysis
Ours
BEVFusion [36]
Camera Image
Ours
BEVFusion [36]
Camera+lidar
Camera only
@@
Figure 5: (a, b) Visualisation of where camera features of ground-truth objects are projected onto the BEV grid for our method compared to BEVFusion [36]. We observe that our method is able to place camera features around objects more narrowly than BEVFusion, which is based on monocular depth estimation. (c) Comparison of saliency maps, cropped to aid visualisation, given the camera image (top) for models trained with camera-lidar (middle) and camera only (bottom). When trained with both camera and lidar, our model selects camera features in an area that is different than when trained with camera only, while [36] behaves similarly in both settings.
@@
We visualise where camera features are projected onto the BEV grid and compare our method to BEVFusion [36]. For our method, we examine the attention map of the final cross-attention block in the transformer, averaged over all attention heads. For BEVFusion, we use the the monocular depth estimate to establish the strength of correspondence between positions in camera and BEV space. We consider only the pixels corresponding to ground-truth objects when calculating the total weight of projected camera features in BEV, see Sec. B.2 for details. As can be seen in Fig. 5(a), our method places camera features predominantly in regions where ground-truth bounding boxes are present. This indicates that it can effectively leverage the lidar point cloud as a context to project camera features at the relevant location in BEV. Compared to BEVFusion shown in Fig. 5(b), the distribution of features appears more narrowly localised and stronger around objects. This could be because our projection mechanism does not require the weights of the camera features to be normalised along their ray, giving our model more flexibility to place features at the desired location. Note that, even though our method also projects camera features outside of ground-truth boxes in BEV, the strength of the activation in these regions is suppressed by the fusion module, see Fig. S4. This is consistent with our findings in Sec. 3, where we showed that the latter part of the model can effectively suppress camera features at the wrong location. More examples can be found in Sec. B.2.
@@
We further explore which pixels in the camera images are most attended to using saliency maps [49]. These are derived by computing the gradient of the maximum class logit with respect to a camera image $I_{j}$, given object query index $i$ and probabilities $z$, as $\partial z_{i,\hat{c}}/\partial I|_{I_{j}}$ where $\hat{c}=\arg\max_{c}z_{i,c}$. They allow us to visualise the contribution of individual pixels to the final detection for a selected object, see Fig. 5(c). Interestingly, we observe that when trained with both camera and lidar, our model tends to select camera features at different locations than when trained with cameras only. In the absence of lidar, our method selects camera features across the entirety of the object, while in the presence of both lidar and cameras, the model selects camera features mainly from the upper part of the object. We observe that this pattern is mostly prevalent for objects close to the ego which are well-represented by lidar point clouds but fades away for far-away objects or objects with few lidar points such as pedestrians, see Fig. S2 for more examples. We hypothesise that our projection architecture enables the model to select camera features that best complement the information encoded in lidar, resulting in differing attention patterns between camera-only and fusion settings. This pattern is less present in BEVFusion [36], which attends to the broader neighbourhood of pixels surrounding the selected object in both cases.
@@ 5.3 Temporal feature aggregation
We train TFA models on sequences of 3 frames for 10 epochs starting from our single frame model’s object detection head, lidar and camera backbones. During training, the lidar and camera backbones are frozen. For inference, we accumulate BEV features for the entire length of a run, yielding detections at each time step. Table 1 shows that temporal feature aggregation boosts object detection performance significantly in all configurations.
@@ 5.4 Ablation experiments
Fusion module Cat+Conv 70.43 71.9
@@
# decoder blocks 1 block 70.29 71.9
@@
Table 2: Impact of model modifications on 3D object detection performance: (i) feature fusion module, (ii) number of transformer decoder blocks in the “Attend” stage, (iii) number of frames in Temporal Feature Aggregation (TFA).  frozen camera backbone.
We ablate some design choices for our method and show their impact on object detection performance on Tab. 2. All ablation experiments use a simpler training schedule with 10 epochs, batch accumulation instead of full batch training and no camera augmentations. We first analyse the effect of different fusion modules: we compare a simple skip connection (add), a small concatenation and convolution layer (Cat+Conv as in [36]) and a gated sigmoid block [31]. We find all perform similarly, with Cat+Conv achieving the best mAP, contrary to findings of [31]. We also ablate the number of transformer decoder blocks in the “Attend” stage of our projection and show that increasing their number leads to a small improvement in mAP. This suggests that our method’s performance scales with compute. We use a single decoder block in our main experiments to balance quality and performance. Finally, we see good improvement in NDS with an increasing number of frames in TFA during training.
@@ 6 Conclusion
In this work, we analysed the role of monocular depth prediction in recent state-of-the-art camera-lidar fusion methods and showed that, surprisingly, improvements in depth estimation did not lead to better object detection performance. Strikingly, we also showed that removing depth estimation altogether did not worsen performance significantly. This led us to hypothesise that relying on monocular depth estimation could be an unnecessary architectural bottleneck when fusing camera and lidar, and prompted us to introduce a novel fusion method that directly combines camera and lidar features using a simple attention mechanism. Compared to projecting camera features using monocular depth, our method allows camera features to contribute to multiple locations in BEV space and gives our model greater flexibility to select complementary camera and lidar features. Finally, we validated the effectiveness of our method on the nuScenes dataset and showed that it improves object detection performance over baselines based on monocular depth estimation and showcased the role of attention as a key contributor to these improvements. We leave detailed investigation of our model in the camera-only setting and inclusion of radar as future work. We hope that our findings will motivate discussions around the role of monocular depth prediction in camera-lidar fusion and prompt further developments in multimodal perception.
@@ Acknowledgements
We thank our FiveAI and Bosch colleagues, especially Tom Joy and Anthony Knittel, for their valuable feedback. We are grateful to Bhavesh Garg for his initial experiments into depth supervision and to Blaine Rogers for initiating work on multimodal fusion at FiveAI.
@@ A.4 Extended depth quality results
We evaluate the performance of the depth classifier using five of the metrics proposed in [10]: root mean squared error (RMSE), root mean squared logarithmic error (RMSLE), mean absolute relative error (Abs. Rel.), mean squared relative error (Sq. Rel.) and fraction outside 125% (Frac. 125). We show all the metrics for 2 different methods of translating the classification output into a depth map: “mode” — in which we use the bin with maximum probability, and “mean” — where the predicted depth is the weighted average of all the bins. We take the average of these quantities over all predictions made by the depth classifier for which we have ground truth to compute the metrics. Camera feature cells for which lidar depth is unspecified are ignored. We compare BEVFusion and four different variants: adding depth supervision using Eq. 2 with various weights $\lambda$, using lidar depth maps instead of monocular depth estimation (lidar), using a pretrained and frozen depth classifier (pretrained), and finally removing depth estimation altogether by projecting camera features at all depths uniformly using Eq. 3 (uniform depth). Quantitative results can be seen in Tab. S1 and qualitative comparisons in Fig. S1.
@@ B Detailed experimental results
Table S2: Per-class object detection scores on the nuScenes validation set (top) and test set (bottom). TFA: Temporal Feature Aggregation. indicates ensembling + TTA.
@@ B.2 Detailed qualitative results
Ours
BEVFusion [36]
Ours
BEVFusion [36]
Cam+lidar
Cam-only
Ours
BEVFusion [36]
Ours
BEVFusion [36]
Cam+lidar
Cam-only